# Deep Multimodal Learning with Missing Modality: A Survey

### v3

@@ Abstract.
During multimodal model training and testing, certain data modalities may be absent due to sensor limitations, cost constraints, privacy concerns, or data loss, negatively affecting performance. Multimodal learning techniques designed to handle missing modalities can mitigate this by ensuring model robustness even when some modalities are unavailable. This survey reviews recent progress in Multimodal Learning with Missing Modality (MLMM), focusing on deep learning methods. It provides the first comprehensive survey that covers the motivation and distinctions between MLMM and standard multimodal learning setups, followed by a detailed analysis of current methods, applications, and datasets, concluding with challenges and future directions.
multimodal learning, deep learning, representation learning, multimodal fusion, audio signal processing, multimedia, missing modality, imperfect modality, incomplete modality
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
1. Introduction
(a) Trends in Missing Modality Publications (2014-2024 October)
(b) Full-Modality Samples vs. Missing-Modality Samples
Figure 1. (a) The trend of papers published on deep multimodal learning with missing modality in the past 10 years. The number of publications has increased over time and has received widespread attention from the community. (b) Description of a three-modality scenario with full- and missing-modality samples. We abbreviate “Modality” as “Mod” in all figures of this paper and use dashed boxes with fading colors to represent missing modalities/modules.
@@ 1. Introduction
Multimodal learning has become a crucial field in Artificial Intelligence (AI). It focuses on jointly analyzing various data modalities, including visual, textual, auditory, and sensory information. This approach mirrors the human capacity to combine multiple senses for better understanding and interaction with the environment. Modern multimodal models leverage the robust generalization capabilities of deep learning to uncover complex patterns and relationships that single-modal systems might not detect. This capability is advancing work across multiple domains (Bayoudh et al., 2022; Hu et al., 2023; Streli et al., 2023). Recent multimodal learning surveys highlight the significant impact of multimodal approaches, demonstrating their ability to enhance performance and enable more sophisticated AI applications (Baltrušaitis et al., 2018; Zhao et al., 2024).
@@
However, real-world multimodal systems often face the challenge of having to handle cases where certain data modalities may be missing or incomplete. This occurs due to various factors such as sensor malfunctions, hardware limitations, privacy concerns, environmental interference, and data transmission issues. In recent years, more and more researchers have shown great interest in this field, which can be demonstrated by the bar chart in Figure 1(a). As illustrated in  Figure 1(b), in a three-modality scenario, data samples can be categorized as either full-modality (containing information from all three modalities) or missing-modality (lacking data from one or more modalities). These problems can arise at any stage from data collection to deployment, significantly impacting model performance. For instance, early affective computing researchers (Cohen et al., 2004; Sebe et al., 2005) found that when collecting facial and audio data, some image or audio samples were not useful due to excessive microphone noise or camera obstructions. This has forced them to propose audio-visual models that can handle the missing modality to recognize human emotional states. In the field of medical AI, privacy concerns and the challenges of obtaining certain data modalities during surgeries or invasive treatments often lead to missing modalities in multimodal datasets (Zhang et al., 2023c). Similarly, in space exploration, NASA’s Ingenuity Mars helicopter (Donaldson, 2024) faced a missing modality challenge when its inclinometer failed due to extreme temperature cycles on Mars. To address this issue, NASA applied a software patch that modified the initialization of navigation algorithms (Team, 2022). Moreover, the inherent structural or quantitative heterogeneity of sensors across different versions or brands of equipment can lead to variations in available modalities. This requires models to be capable of handling inputs with differing types of modalities in real-world applications. The unpredictability of real-world scenarios and diversity of data sources further compound this challenge. Consequently, developing robust multimodal systems that can perform effectively with missing modalities has become a critical focus in the field.
Removing missing-modality samples is a common and simple multimodal dataset pre-processing strategy when the dataset has some missing-modality samples. But this method will waste the valuable information contained in missing-modality samples and cannot help models handle missing modality samples during testing, as it can only be used as a “temporary” solution for training. This has prompted many researchers to propose many multimodal learning methods for missing-modality smap. In this survey, we refer to the challenge of handling missing modalities for multimodal learning as the “missing modality problem.” We name solutions to this problem as Multimodal Learning with Missing Modality (MLMM). Those approaches contrast with the conventional setup that utilizes the full set of modalities, which we name as Multimodal Learning with Full Modality (MLFM). Specifically, in MLFM tasks, given an $\mathbf{N}$-modality dataset, we are typically required to train and test a model that can process and fuse all $\mathbf{N}$ modalities to make decisions. In contrast, MLMM tasks may use fewer than the $\mathbf{N}$ modalities available during training or testing due to factors such as data collection limitations or constraints in deployment environments. The primary challenge in MLMM lies in dynamically and robustly handling and fusing information from any number of available modalities during training and testing, yet maintaining performance comparable to that achieved with full-modality samples.
Figure 2. Our taxonomy of deep multimodal learning with missing modality methods. We categorize existing methods into two aspects: data processing and strategy design. Data Processing: we differentiate between modality imputation (handling at the modality data level) and representation-focused models (dealing with at the data representation level). Strategy Design: we distinguish between architecture-focused models (model architecture adjustments) and model combinations (combining multiple models externally). “MLLMs”: multimodal large language models.
@@
This survey encompasses recent advances in MLMM and its applications across various domains, including information retrieval (Malitesta et al., 2024), remote sensing (Wei et al., 2023), robotic vision (Gunasekar et al., 2020), medical diagnosis (Zhou et al., 2023; Shah et al., 2023; Azad et al., 2022), sentiment analysis (Wagner et al., 2011), and multi-view clustering (Chao et al., 2021). We also introduce a fine-grained taxonomy of MLMM methodologies, application scenarios, and corresponding datasets. The main contributions of this survey are: (1) A comprehensive survey of MLMM methods across diverse domains, accompanied by an extensive compilation of relevant datasets, highlighting the versatility of MLMM in addressing real-world challenges. (2) A novel, fine-grained taxonomy of MLMM methodologies, with a multi-faceted categorization framework based on multi-modal integration stages and missing modality recovery strategies. (3) An in-depth analysis of current MLMM approaches, their challenges, and future research directions, contextualized within the proposed taxonomical framework.
@@
Paper Collection: In our literature search methodology, we primarily source papers from Google Scholar and major conferences/journals in AI, Machine Learning, Computer Vision, Natural Language Processing, Audio Signal Processing, Data Mining, Multimedia, Medical Imaging, and Remote Sensing. The collected papers are from, but not limited to, top-tier conferences (e.g., AAAI, IJCAI, NeurIPS, ICLR, ICML, CVPR, ICCV, ECCV, ACL, EMNLP, KDD, ACM MM, MICCAI, ICASSP) and journals (e.g., TPAMI, TIP, TMI, TMM, JMLR). Please refer to Section A in the Supplementary Materials for the full names of these conferences and journals. We have compiled a total of 315 significant papers from the period spanning 2012 to October 2024. Our search strategy involved using keywords such as “incomplete,” “missing,” “partial,” “absent,” and “imperfect,” combined with terms like “multimodal learning,” “deep learning,” “representation learning,” “multi-view learning,” and “neural networks.”
@@
Survey Organization: Firstly, we explain the background and motivation of this survey in Section 1. In Section 2, we introduce our taxonomy and categorize existing deep MLMM methods from a methodological perspective, detailing them in two aspects and four types (Figure 2). In the following Section 3 and Section 4, we introduce various methods from the aspects of model data processing and strategy design respectively. We then summarize current application scenarios and corresponding datasets used in this field in Section 6. In Section 7, we discuss unresolved challenges and future directions. Finally, we present our conclusions drawn from the exploration of deep MLMM in Section 8.
2. Methodology Taxonomy: An Overview
@@
We review current deep MLMM methods from two key aspects: data processing and strategy design.
2.1. Data Processing Aspect
@@
Methods that focus on exploring the data processing aspect of the model can be divided into Modality Imputation and Representation-Focused Models, according to whether the processing of missing modality occurs at the modality data level or the data representation level.
@@
(1) Modality Imputation operates at the modality data level and fills in the missing information by compositing (Chen et al., 2020; Sun et al., 2024a; Parthasarathy and Sundaram, 2020; Yang et al., 2024, 2024; Campos et al., 2015) (modality composition methods) or generating (Chen et al., 2024b; Ma et al., 2021b; Arya and Saha, 2021; Wang et al., 2024a; Tran et al., 2017; Zhang et al., 2023b) absent modalities (modality generation methods) from available modalities. Those approaches are rooted in the idea that if a missing modality can be accurately imputed, downstream tasks can continue as if “full” modalities were available.
@@
(2) Representation-Focused Models are designed to address missing modalities at the representation level. In some cases, the coordinated representation methods (Liang et al., 2019; Wang et al., 2020a; Ma et al., 2021c) impose some specific constraints on the representations of different modalities to help align the representations of different modalities in the semantic space, so that the model can be effectively trained even when facing missing modalities. Other representation-focused methods either generate the missing-modality representation using available data (Zheng et al., 2021; Li et al., 2024b; Hoffman et al., 2016; Ma et al., 2021b; Park et al., 2024) or combine the representations of existing modalities (Zhi et al., 2024; Sun et al., 2024a; Zhou et al., 2021; Delgado-Escano et al., 2021; Zheng et al., 2024; Zhang et al., 2024) to fill in the gaps.
2.2. Strategy Design Aspect
@@
(1) Architecture-Focused Models address missing modalities by designing flexible model architectures that can adapt to varying numbers of available modalities during training or inference. A key technique here is based on attention mechanisms (Ge et al., 2023; Yao et al., 2024; Gong et al., 2023; Mordacq et al., 2024; Radosavovic et al., 2024; Jang et al., 2024; Qiu et al., 2023), which dynamically adjusts the modality fusion and processing, allowing the model to handle any number of input modalities. Another approach is based on knowledge distillation (Wang et al., 2020b; Saha et al., 2024; Zhang et al., 2023c; Chen et al., 2024a; Wang et al., 2023a; Shi et al., 2024; Wang et al., 2023d), where the model is trained to accommodate missing modalities by transferring knowledge from full-modality models to those operating with incomplete data or distilling between different branches inside the model. Additionally, graph learning-based methods (Zhao et al., 2022; Zhang et al., 2022; Lian et al., 2023; Malitesta et al., 2024) exploit the natural relationships between modalities, using graphs to dynamically fuse and process available modalities while compensating for missing ones. Finally, MLLMs (Zhan et al., 2024; Wu et al., 2023a; Li et al., 2023c) also play a crucial role in this category, as their ability to handle long contexts and act as feature processors enables them to accept and process representations from any number of modalities. These architectural strategies collectively allow models to maintain performance even when dealing with incomplete multimodal inputs.
(2) Model Combinations tackle missing modality problems by employing strategies that leverage multiple models or specialized training techniques. One approach is to use dedicated training strategies (Zeng et al., 2022; Chen et al., 2022; Xue and Marculescu, 2023) tailored for different modality cases, ensuring that each case is trained for optimal performance. Another approach involves ensemble methods (Hu et al., 2020; Wang et al., 2021; Lee et al., 2023b), where models trained on either partial/full sets of modalities are combined, allowing the system to select the most suitable model based on the available modalities to do joint predictions. Additionally, discrete scheduler methods (Wu et al., 2023b; Shen et al., 2024; Surís et al., 2023) can incorporate various downstream modules to flexibly process any number of modalities and handle specific tasks. These schedulers intelligently select and combine the outputs of multiple models or modules to manage missing-modality scenarios, offering a versatile solution for multimodal tasks.
@@
3. Methodologies in Data Processing Aspect
3.1. Modality Imputation
@@
Figure 3. Zero/Random values composition methods. If we assume modality 2 is missing, then this modality will be replaced with zero/random values. “DNN” in all figures of this survey means different kinds of deep neural networks.
@@
Figure 4. Retrieval-based modality composition methods search for one or more samples by randomly selecting or using simple retrieval algorithms like KNN, or its variants, from same-category samples that have the required missing modalities, and then compose them with the input missing-modality sample to form a “full”-modality sample.
3.1.1. Modality Composition Methods
@@
Modality composition methods are widely employed for its simplicity and effectiveness to maintain the original dataset size. Zero/Random Values Composition Methods represent a type of modality composition method that replaces a missing modality with zeros or random values, as shown in Figure 3. In recent research (Chen et al., 2020; Sun et al., 2024a; Liu et al., 2023c; Malitesta et al., 2024), they are often used as a baseline method for comparison with other more sophisticated methods. For the missing sequential data problem, such as missing frames in videos, the similar Frame-Zero method (Parthasarathy and Sundaram, 2020) was proposed to replace the missing frames. These methods are prevalent in typical multimodal learning procedures and can be used to balance and integrate information from different modalities when making predictions. This prevents the model from over-relying on dominant modalities available for each sample, enhancing its robustness by encouraging a more balanced integration of information across all available modalities.
@@
Retrieval-Based Representation Composition Methods (Figure 4) represent another type of modality composition method, which replaces the missing modality data by copying or averaging data from retrieved samples with the same classification. Some other methods randomly select a sample that has the same classification and required missing modality from other samples. The selected modality data is then composed with the missing-modality sample to form a full-modality sample for training. But those retrieval-based modality composition methods are not applicable to pixel-level tasks, such as segmentation, and are only suitable for simple tasks (e.g., classification) because they may lead to the overfitting of noisy data, if mismatched samples are combined. For example, Yang et al. (Yang et al., 2024) proposed Modal-mixup to complete the training datasets by randomly complementing same-category samples with missing modalities. However, such methods cannot solve the missing modality problem during the testing phase because they rely on known labels of training data samples. In some multimodal streaming data classification tasks, such as audio-visual expression recognition, video streams may experience frame drops due to network communication packet loss, etc. Frame-Repeat (Parthasarathy and Sundaram, 2020) was proposed to to make up for the missing frames by using past frames.
@@
Other methods (Yang et al., 2024; Campos et al., 2015) also used K-Nearest Neighbors (KNN), or its variants, to retrieve the best-matched samples for composition. For these matched samples, they select the sample with the highest score or obtain the average values of these samples to supplement the missing modality data. Their experiments have shown that KNN-based methods generally perform better than the above methods, and can handle missing modality during testing. Nevertheless, such KNN retrieval-based modality composition methods often suffer from high computational complexity, sensitivity to imbalanced data, and significant memory overhead.
@@
Figure 5. Description of two typical modality generation methods. We set modality 2 as the missing modality for examples and use other available modalities to generate modality 2. “GEN” in both figures represents modality generation networks. (a) We set up a modality-2 generator (GEN-2) from other modalities. (b) All modalities are input and generated together by using a single GEN.
3.1.2. Modality Generation Methods

### v4

@@ Abstract
During multimodal model training and testing, certain data modalities may be absent due to sensor limitations, cost constraints, privacy concerns, or data loss, which can degrade performance. Multimodal learning techniques that explicitly account for missing modalities aim to improve robustness by enabling models to perform reliably even when certain inputs are unavailable. This survey presents the first comprehensive review of Multimodal Learning with Missing Modality (MLMM), with a focus on deep learning approaches. We outline the motivations and key distinctions between MLMM and conventional multimodal learning, provide a detailed analysis of existing methods, applications, and datasets, and conclude by highlighting open challenges and future research directions.
1 Introduction
@@
Multimodal learning has become a crucial field in Artificial Intelligence (AI). It focuses on jointly analyzing various data modalities, including visual, textual, auditory, and sensory information. This approach mirrors the human capacity to combine multiple senses for better understanding and interaction with the environment. Modern multimodal models leverage the robust generalization capabilities of deep learning to uncover complex patterns and relationships that single-modal systems might not detect. This capability is advancing work across multiple domains (Bayoudh et al., 2022; Hu et al., 2023; Streli et al., 2023). Recent surveys show that multimodal learning significantly boosts performance and enables more advanced AI applications (Baltrušaitis et al., 2018; Zhao et al., 2024; Wu et al., 2024a).
However, real-world multimodal systems frequently encounter scenarios where certain data modalities are missing or incomplete. Such cases arise from sensor malfunctions, hardware limitations, privacy concerns, environmental interference, and data transmission failures. Interest in addressing this challenge has grown considerably in recent years, as reflected in the publication trends shown in figure 1(a). For instance, in a three-modality setting, data samples can be classified as either full-modality (containing all three modalities) or missing-modality (lacking one or more). Missing modalities can occur at any stage, from data collection to deployment, and can substantially degrade model performance.
@@
Early works in affective computing (Cohen et al., 2004; Sebe et al., 2005) reported that when collecting facial and audio data, some image or audio samples were not useful due to excessive microphone noise or camera obstructions. This has forced them to propose audio-visual models that can handle the missing modality to recognize human emotional states. In the field of medical AI, privacy concerns and the challenges of obtaining certain data modalities during surgeries or invasive treatments often lead to missing modalities in multimodal datasets (Zhang et al., 2023b; Li et al., 2025b). Similarly, in space exploration, NASA’s Ingenuity Mars helicopter (Donaldson, 2024) faced a missing modality challenge when its inclinometer failed due to extreme temperature cycles on Mars. To address this issue, NASA applied a software patch that modified the initialization of navigation algorithms (Team, 2022). Beyond such domain-specific cases, structural or quantitative heterogeneity across sensor types, versions, or brands can also yield inconsistent modality availability. Combined with the unpredictability of real-world scenarios and the diversity of data sources, these factors make robustness to missing modalities an essential requirement for multimodal systems. Consequently, developing models that can effectively handle incomplete inputs has become a central research focus.
In this survey, we focus on the challenge of handling incomplete inputs in multimodal learning, which we refer to as the missing modality problem. We term solutions to this challenge Multimodal Learning with Missing Modality (MLMM). Unlike the conventional full-modality setting, where all $\mathbf{N}$ modalities are assumed to be consistently available, MLMM must operate when one or more modalities are absent during training or testing. The central goal of MLMM is to robustly exploit the available modalities while maintaining performance comparable to systems trained with complete modality information.
@@
This survey reviews recent advances in MLMM and its applications across diverse domains, including information retrieval (Malitesta et al., 2024; Pipoli et al., 2025), remote sensing (Wei et al., 2023), Pervasive Computing (Nweke et al., 2018; Van Der Donckt et al., 2024; Chen et al., 2023a), robotic vision (Gunasekar et al., 2020; Park et al., 2025), (bio-)medical domain (Zhou et al., 2023; Shah et al., 2023; Azad et al., 2022; Zhu et al., 2025), sentiment analysis (Wagner et al., 2011; Wang et al., 2025b), and multi-view clustering (Chao et al., 2021). We also introduce a fine-grained taxonomy of MLMM methodologies, application scenarios, and corresponding datasets.
(a) Trends in Missing Modality Publications
(b) Full-Modality vs. Missing-Modality Samples
Figure 1: (a) The trend of papers published on deep multimodal learning with missing modality in the past 10 years (2014-2025.8). The number of publications has increased over time and has received widespread attention from the community. (b) Description of a three-modality scenario with full- and missing-modality samples. We abbreviate “Modality” as “Mod” in all figures of this paper and use dashed boxes with fading colors to represent missing modalities/modules.
Figure 2: Our taxonomy of deep multimodal learning with missing modality methods. We categorize existing methods into two aspects: data processing and strategy design. Data Processing: we differentiate between modality imputation (handling at the modality data level) and representation-focused models (dealing with at the data representation level). Strategy Design: we distinguish between architecture-focused models (model architecture adjustments) and model combinations (combining multiple models externally). “MLLMs”: multimodal large language models.
Contributions: (1) A comprehensive survey of MLMM methods across diverse domains, accompanied by an extensive compilation of relevant datasets, highlighting the versatility of MLMM in addressing real-world challenges. (2) A novel, fine-grained taxonomy of MLMM methodologies, with a multi-faceted categorization framework based on multi-modal integration stages and missing modality recovery strategies. (3) An in-depth analysis of current MLMM approaches, their challenges, and future research directions, contextualized within the proposed taxonomical framework.
@@
Paper Collection: In our literature search methodology, we primarily source papers from Google Scholar and major conferences/journals in AI, Machine Learning, Computer Vision, Natural Language Processing, Audio Signal Processing, Data Mining, Multimedia, Medical Imaging, Remote Sensing and Pervasive Computing. The collected papers are from, but not limited to, top-tier conferences (e.g., AAAI, IJCAI, NeurIPS, ICLR, ICML, CVPR, ICCV, ECCV, ACL, EMNLP, KDD, ACM MM, MICCAI, ICASSP) and journals (e.g., TPAMI, TIP, TMI, TMM, JMLR, TMLR). Please refer to Section A in the Supplementary Materials for the full names of these conferences and journals. We have compiled a total of 354 significant papers from the period spanning 2012 to August 2025. Our search strategy involved using keywords such as “incomplete,” “missing,” “partial,” “absent,” and “imperfect,” combined with terms like “multimodal learning,” “deep learning,” “representation learning,” “multi-view learning,” and “neural networks.”
Survey Organization: Firstly, we explain the background and motivation of this survey in section 1.
@@
In section 2, we introduce our taxonomy and categorize existing deep MLMM methods from a methodological perspective, detailing them in two aspects and four types (figure 2). In the following section 3 and section 4, we introduce various methods from the aspects of model data processing and strategy design respectively. We then summarize current application scenarios and corresponding datasets used in this field in section 6. In section 7, we discuss unresolved challenges and future directions. Finally, we present our conclusions drawn from the exploration of deep MLMM in section 8.
2 Methodology Taxonomy: An Overview
@@
We review current deep MLMM methods from two key aspects: data processing and strategy design, based on the contributions of existing work.
2.1 Data Processing Aspect
@@
Methods that focus on exploring the data processing aspect of a model or framework can be divided into Modality Imputation and Representation-Focused Models, depending on whether the model’s handling of missing modalities occurs at the modality data level or at the data representation level.
@@
(1) Modality Imputation operates at the modality data level and fills in the missing information by compositing (Chen et al., 2020; Sun et al., 2024b; Parthasarathy and Sundaram, 2020; Yang et al., 2024; Campos et al., 2015) (modality composition methods) or generating (Chen et al., 2024a; Ma et al., 2021c; Arya and Saha, 2021; Wang et al., 2024c; Tran et al., 2017; Zhang et al., 2023c) absent modalities (modality generation methods) from available modalities. Those approaches are rooted in the idea that if a missing modality can be accurately imputed, downstream tasks can continue as if “full” modalities were available.
@@
(2) Representation-Focused Models are designed to address missing modalities at the representation level. In some cases, the coordinated representation methods (Liang et al., 2019; Wang et al., 2020b; Ma et al., 2021b) impose some specific constraints to help align the representations of available modalities in the same semantic space, so that the model can be effectively trained even when facing missing modalities. Other representation-focused methods either generate the missing-modality representation using available data (Zheng et al., 2021; Li et al., 2024b; Hoffman et al., 2016; Ma et al., 2021c; Park et al., 2024) or combine the representations of existing modalities (Zhi et al., 2024; Sun et al., 2024b; Zhou et al., 2021; Delgado-Escano et al., 2021; Zheng et al., 2024; Zhang et al., 2024) to fill in the gaps.
2.2 Strategy Design Aspect
@@
(1) Architecture-Focused Models address missing modalities by designing flexible model architectures that can adapt to varying numbers of available modalities during training or inference. A key technique here is based on attention mechanisms (Ge et al., 2023; Yao et al., 2024; Gong et al., 2023; Mordacq et al., 2024; Radosavovic et al., 2024; Jang et al., 2024; Qiu et al., 2023), which dynamically adjusts the modality fusion and processing, allowing the model to handle any number of input modalities. Another approach is based on knowledge distillation (Wang et al., 2020a; Saha et al., 2024; Zhang et al., 2023b; Chen et al., 2024b; Wang et al., 2023a; Shi et al., 2024; Wang et al., 2023c), where the model is trained to accommodate missing modalities by transferring knowledge from full-modality models to those operating with incomplete data or distilling between different branches inside the model. Additionally, graph learning-based methods (Zhao et al., 2022; Zhang et al., 2022; Lian et al., 2023; Malitesta et al., 2024) exploit the natural relationships between modalities, using graphs to dynamically fuse and process available modalities while compensating for missing ones. Finally, MLLMs (Zhan et al., 2024; Wu et al., 2023b; Li et al., 2023a) also play a crucial role in this category, as their ability to handle long contexts and act as feature processors enables them to accept and process representations from any number of modalities. These architectural strategies collectively allow models to maintain performance even when dealing with incomplete multimodal inputs.
(2) Model Combinations tackle missing modality problems by employing strategies that leverage multiple models or specialized training techniques. One approach is to use dedicated training strategies (Zeng et al., 2022; Chen et al., 2022; Xue and Marculescu, 2023) tailored for different modality cases, ensuring that each case is trained for optimal performance. Another approach involves ensemble methods (Hu et al., 2020; Wang et al., 2021; Lee et al., 2023b), where models trained on either partial/full sets of modalities are combined, allowing the system to select the most suitable model based on the available modalities to do joint predictions. Additionally, discrete scheduler methods (Wu et al., 2023a; Shen et al., 2024; Surís et al., 2023) can incorporate various downstream modules to flexibly process any number of modalities and handle specific tasks. These schedulers intelligently select and combine the outputs of multiple models or modules to manage missing-modality scenarios, offering a versatile solution for multimodal tasks.
@@
3 Methodologies in Data Processing Aspect
3.1 Modality Imputation
@@
Figure 3: Zero/Random values composition methods. If we assume modality 2 is missing, then this modality will be replaced with zero/random values. “DNN” in all figures of this survey means different kinds of deep neural networks.
@@
Figure 4: Retrieval-based modality composition methods search for one or more samples by randomly selecting or using simple retrieval algorithms like KNN, or its variants, from same-category samples that have the required missing modalities, and then compose them with the input missing-modality sample to form a “full”-modality sample.
3.1.1 Modality Composition Methods
@@
Modality composition methods are widely employed for its simplicity and effectiveness to maintain the original dataset size. Zero/Random Values Composition Methods represent a type of modality composition method that replaces a missing modality with zeros or random values, as shown in figure 3. In recent research (Chen et al., 2020; Sun et al., 2024b; Liu et al., 2023a; Malitesta et al., 2024), they are often used as a baseline method for comparison with other more sophisticated methods. For the missing sequential data problem, such as missing frames in videos, the similar Frame-Zero method (Parthasarathy and Sundaram, 2020) was proposed to replace the missing frames.
These methods are prevalent in typical multimodal learning procedures and can be used to balance and integrate information from different modalities when making predictions. This prevents the model from over-relying on dominant modalities available for each sample, enhancing its robustness by encouraging a more balanced integration of information across all available modalities.
@@
Retrieval-Based Representation Composition Methods (figure 4) represent another type of modality composition method, which replaces the missing modality data by copying or averaging data from retrieved samples with the same classification. Some other methods randomly select a sample that has the same classification and required missing modality from other samples. The selected modality data is then composed with the missing-modality sample to form a full-modality sample for training. But those retrieval-based modality composition methods are not applicable to pixel-level tasks, such as segmentation, and are only suitable for simple tasks (e.g., classification) because they may lead to the overfitting of noisy data, if mismatched samples are combined. For example, Yang et al. (2024) proposed Modal-mixup to complete the training datasets by randomly complementing same-category samples with missing modalities. However, such methods cannot solve the missing modality problem during the testing phase because they rely on known labels of training data samples. In some multimodal streaming data classification tasks, such as audio-visual expression recognition, video streams may experience frame drops due to network communication packet loss, etc. Frame-Repeat (Parthasarathy and Sundaram, 2020) was proposed to to make up for the missing frames by using past frames.
@@
Other methods (Yang et al., 2024; Campos et al., 2015) also used K-Nearest Neighbors (KNN), or its variants, to retrieve the best-matched samples for composition. For these matched samples, they select the sample with the highest score or obtain the average values of these samples to supplement the missing modality data. Their experiments have shown that KNN-based methods generally perform better than the above methods, and can handle missing modality during testing. Nevertheless, such KNN retrieval-based modality composition methods often suffer from high computational complexity, sensitivity to imbalanced data, and significant memory overhead. In pervasive computing, some researcher (Hussein et al., 2024a) proposed to cluster sensor data and pre-learns the best imputation pattern for each cluster combination, enabling fast lookup-based completion of missing modality at runtime.
@@
Figure 5: Description of two typical modality generation methods. We set modality 2 as the missing modality for examples and use other available modalities to generate modality 2. “GEN” in both figures represents modality generation networks. (a) We set up a modality-2 generator (GEN-2) from other modalities. (b) All modalities are input and generated together by using a single GEN.
3.1.2 Modality Generation Methods